# Independence Assumption RLS Regularized Least Squares

### v1

@@ 4.1 A computable lower bound on epistemic uncertainty
$I(\mu)=\tilde{\Omega}\Big{(}(1-\mathbb{E}[U_{k}])\,\widehat{I}_{k}(\gamma)-\mathbb{E}[U_{k}]\Big{)}\qquad k\to\infty~.$
@@
Despite these pessimistic bounds, in reality we expect the expected missing mass to be significantly smaller, especially when $\mu$ is heavy-tailed. It is well-known that natural languages (and many artificial ones) follow a Zipf distribution, where probability of each word (or a text piece) is proportional to $1/\mathrm{freq(text)}^{\alpha}$ for some exponent $\alpha>1$, where $\mathrm{freq}()$ is a frequency of occurrence in the corpus (Piantadosi, 2014). Then, we expect that $\mathbb{E}[U_{k}]$ should be much smaller than in such a case, since sampling from the tail of Zipf distribution is a rare event. To this end, in Appendix˜C we show that if $\widetilde{Q}$ is Zipf with exponent $\alpha>1$, then for any free parameter $\beta>0$,
@@
Finally in Section˜C.4 we present a data-dependent estimation of $\mathbb{E}[U_{k}]$ based on a concentration inequality for a missing mass and repetitive sampling from LLM, in the context of some Q/A datasets. We conclude that the expected missing mass is very small: Most of the upper bounds on the expected missing mass (we have one upper bound per question) are highly concentrated close to 0.
@@ 5 Score-based hallucination tests
Let $\widehat{I}_{k}(\gamma,x)\equiv\widehat{I}_{k}(\gamma)$ computed as in Algorithm˜1 for $\mu=\widetilde{Q}$, to emphasize the explicit dependence on the query $x$. The uncertainty estimate $\widehat{I}_{k}(\gamma,x)$ derived above can be used as a score indicating the strength of our belief that the LLM hallucinates for the given query $x$. Such a score can then be used to design abstention policies: if the response is deemed to be hallucinated, the system abstains from responding, while a response is provided otherwise. Score-based abstention methods usually compute a score chosen by the user (such as the response likelihood or the estimator $\widehat{I}(\gamma)$ discussed earlier), and declare hallucination if the score is above or below a threshold, which is determined through calibration. To detect hallucinations successfully, the threshold can be adjusted through calibration on a given task using a hold-out (ground-truth) sample, see, for instance, the paper of Yadkori et al. (2024) where this calibration is discussed in detail.
@@
$a_{\lambda}(x)=\begin{cases}0,&\text{ if }\ \widehat{I}_{k}(\gamma,x)<\lambda;\\ 1,&\text{ if }\ \widehat{I}_{k}(\gamma,x)\geq\lambda;\end{cases}$
@@ 6 Experiments
In this section we evaluate our abstention method derived based on the MI estimate in Section˜5 on a variety of closed-book open-domain question-answering tasks.
@@
Datasets. We consider three different datasets and their combinations: As base datasets, we consider (i) a random subset of 50,00050,00050 , 000 datatpoints from the TriviaQA dataset (Joshi et al., 2017), and (ii) the entire AmbigQA dataset (with 120381203812038 datapoints) (Min et al., 2020). These datasets mostly contain single-label queries, and only contain a few multi-label ones. [^8] Note that the multi-label queries in these datasets typically behave as single-label ones in the sense that the LLM assigns overwhelming probability to a dominant response. Moreover, we created a multi-label dataset based on the WordNet dataset (Fellbaum, 1998): We extracted all (6015) datapoints from WordNet at depth 444 or more of the physical_entity subtree. For each datapoint (entity, children) in WordNet, we constructed a query of the form “Name a type of entity.” and children are considered target labels.
@@
Comparison of responses and computing the output distributions. We use the F1 score999 In this context, the F1 score is calculated based on token inclusion (Joshi et al., 2017, Devlin et al., 2019): for two sequences $a=(a_{1},\ldots,a_{n})$ and $b=(b_{1},\ldots,b_{m})$, defining $p=|a\cap b|/n$ and $r=|a\cap b|/m$ (where $|a\cap b|$ is the size of the intersection of $a$ and $b$, in which for repetitions of an element $y$, we consider the minimum number of repetitions in $a$ and $b$, i.e., $\min_{c\in\{a,b\}}|\{i\,:\,c_{i}=y\}|$, in calculating the size of the intersection) we define $F1=2pr/(p+r)$. Relating to the standard definition of the F1 score, $p$ and $r$ play the role of precision and recall, respectively, if $a$ is thought of as a prediction of $b$. thresholded at 0.250.250.25 to decide if two text sequences match. When multiple responses are sampled, we approximate the output distribution of an LLM in a semantically meaningful way by collapsing matching responses into a single response: we sample $k=10$ responses at temperature 0.90.90.9 for each query, and all those that match (according to the F1 score) are considered identical and their probabilities are aggregated. We only consider queries for which the greedy (temperature zero) and at least one of the random responses are shorter than 202020 characters. This is because the F1 score (as a match function) and log-probabilities (as a measure of uncertainty) are less reliable for longer sequences. After this filtering, we are left with 388703887038870 datapoints for TriviaQA, 531553155315 datapoints for AmbigQA, and 329632963296 datapoints for WordNet.
@@
Baselines. We consider abstention policies based on four scoring methods. The first three are as follows: (i) the probability of the greedy response (denoted by $T0$); (ii) the semantic-entropy method of Kuhn et al. (2023) whose score is the entropy of $k=10$ generated samples (denoted by S.E.). To calculate entropy, we first aggregate probabilities of equivalent responses and normalize the probabilities so that they sum to 1 (as described above); and (iii) our proposed mutual information score as defined in Section˜4 (and denoted by M.I.) with the choices of $k=10$, $n=2$, and $\gamma=0$ (the latter choice approximates the case that the number of potential responses can be very large in which case the theoretical choice of $\gamma$ would be very small). To calculate the mutual information, we first generate $k=10$ random samples. Then for any response $Y$, we calculate the probability of all generated responses given the prompt $F_{1}(x,Y)$. We construct estimates $\widehat{Q}(Y)$ and $\widehat{Q}(Y^{\prime}|Y)$ by aggregating probabilities of equivalent responses, and normalizing the probabilities so that they sum to 1.
@@
In our experiments we either sweep through all abstention thresholds (Figure˜5), or optimize the threshold on some calibration data, as explained in the description of the relevant experiment (Figure˜6).
Results. We consider the precision-recall (PR) trade-off for the various methods on the different datasets. Here, recall is the percentage of queries where the method does not abstain, and precision is the percentage of correct decisions among these queries. [^11] In some figures, for better illustration, we show the error rate which is one minus the precision. Figure˜5ab show PR-curves for the baselines and the proposed method on TriviaQA and AmbigQA. As can be seen, our method is better than the $T0$ and S.V. baselines, but performs similarly to the S.E. method. This is because the TriviaQA and AmbigQA datasets contain mostly single-label queries, and therefore a first-order method such as S.E. is sufficient to detect hallucinations. The AmbigQA dataset contains a few multi-label queries, but upon closer inspection, we observe that the LLM has low entropy on most of these queries. [^12] Such a case can also be seen in the query “Name a city in the UK.” in Figure 3 where the response “London” has probability 0.9580.9580.958. Therefore, a first-order method can perform as well as our method on such queries. Our proposed method, as well as the baselines, make no mistakes on the WordNet dataset (as the prediction of the LLM is always correct), hence we omit those results. The S.V. baseline performs significantly worse than the other methods when the recall is not high (is below about 0.8).
@@
The similar performance for the S.E. and M.I. methods shown in Figure˜5ab is due to the fact that the LLM has low entropy on most multi-label queries. However, ideally, an LLM should have higher entropy on multi-label queries (which would demonstrate broader knowledge, not focusing on a single possible answer). To include such queries, we mix the TriviaQA and AmbigQA datasets with our WordNet-based dataset with “truely” multi-label queries as constructed above. To enhance the intended effect, we filter our WordNet dataset by keeping only queries with entropy higher than 0.70.70.7 (approximately the entropy of the uniform distribution over two atoms). Then we have 842842842 remaining datapoints in WordNet. Note that when considered in isolation, both our proposed method and the semantic entropy method rarely make mistakes on this dataset. Then we create two new datasets by combining our 842842842 WordNet datapoints with 842842842 randomly selected datapoints from TriviaQA and AmbigQA, respectively, resulting in the TriviaQA+WordNet and AmbigQA+WordNet datasets. Figure˜5cd show PR-curves for the S.E. and M.I. methods on these two combined datasets. Apart from low recall values, the performance of the S.E. method degrades noticeably with the addition of extra multi-label data. This precision/recall curve might look somewhat strange (with precision sometimes increasing with recall); this is due to the fact that both methods are always correct on the large number of high-entropy WordNet queries, where the LLM’s default predictions are correct.
@@
The hardness with the combined datasets is that the predominantly single-label datasets (TriviaQA, AmbigQA) might need a different calibration threshold than the multi-label WordNet dataset, and this is better handled by our proposed method than by S.E. To better illustrate the improved abstention properties of our method, we examine how the two methods handle when the output of the LLM is diverse (i.e., has high entropy). In order to do this, we perform the following experiment: We create a calibration dataset by adding 500500500 random datapoints from the WordNet dataset to 500500500 random datapoints from TriviaQA, and another such random dataset for test. We determine the abstention thresholds on the calibration dataset for both the S.E. and the M.E. methods, [^13] This is done by fixing the target loss rates of 0.05 for TriviaQA and 0.15 for AmbigQA, and finding threshold parameters that lead to these rates on the calibration set. and measure the performance (error rate, i.e., 1 minus precision, and recall) of the resulting abstention policies on the test set. We repeat this process 101010 times and report mean values and 95% confidence intervals with Gaussian approximation. We perform a similar evaluation process for mixtures of AmbigQA and WordNet datasets. Figure˜6 show that while the S.E. method has similar recall and error rates to those of the proposed method on low-entropy queries, its recall values are much lower for queries with higher entropy, while the M.E. method makes only few mistakes on these queries.
@@
Figure 6: Recall and error rates (one minus precision: percentage of mistakes when not abstaining) of the proposed and the baseline method on TriviaQA+WordNet and AmbigQA+WordNet datasets. On TriviaQA+WordNet and AmbigQA+WordNet datasets, the methods are calibrated at target loss of 0.050.050.05 and 0.150.150.15, respectively. On the x-axis, the queries are partitioned according to the entropy of the LLM’s output. Error bars show 2 standard deviation confidence intervals (based on 10 repetitions). While the first-order S.E. method has similar recall and error rates to those of the proposed M.E. method on low-entropy queries, its recall values are nearly zero for queries with higher entropy.

### v2

@@ 4.1 A computable lower bound on epistemic uncertainty
$I(\mu)=\tilde{\Omega}\Big{(}(1-\mathbb{E}[U_{k}])\,\widehat{I}_{k}(\gamma_{1},\gamma_{2})-\mathbb{E}[U_{k}]\Big{)}\qquad k\to\infty~{}.$
@@
Despite these pessimistic bounds, in reality we expect the expected missing mass to be significantly smaller, especially when $\mu$ is heavy-tailed. It is well-known that natural languages (and many artificial ones) follow a Zipf distribution, where probability of each word (or a text piece) is proportional to $1/\mathrm{freq(text)}^{\alpha}$ for some exponent $\alpha>1$, where $\mathrm{freq}()$ is a frequency of occurrence in the corpus (Piantadosi, 2014). Then, we expect that $\mathbb{E}[U_{k}]$ should be much smaller than in such a case, since sampling from the tail of Zipf distribution is a rare event. To this end, in Appendix E we show that if $\widetilde{Q}$ is Zipf with exponent $\alpha>1$, then for any free parameter $\beta>0$,
@@
Finally in Section E.4 we present a data-dependent estimation of $\mathbb{E}[U_{k}]$ based on a concentration inequality for a missing mass and repetitive sampling from LLM, in the context of some Q/A datasets. We conclude that the expected missing mass is very small: Most of the upper bounds on the expected missing mass (we have one upper bound per question) are highly concentrated close to 00.
4.2 Taking semantic equivalences into account
Although Theorem 4.5 and Theorem 4.6 provide a lower bound for the divergence between the LLM distribution $Q$ and the ground-truth $P$, the bound might be loose as it ignores the semantic equivalences between texts. Given a semantic equivalence definition, we propose constructing new ground-truth $P^{\prime}$ and LLM distribution $Q^{\prime}$, where the probability of a cluster is the sum of probabilities of all semantically equivalent texts in that cluster. We use a similarity function $s$ to define semantic equivalences: two texts are considered equivalent if their similarity is greater than a given threshold $\tau$. Our choices for similarity functions in the experiments are described in Section 6. We assume the similarity function and the similarity threshold induce a clustering of the space ${\mathcal{X}}$, i.e. $s(Y,Y^{\prime})\geq\tau$ if and only if they are in the same cluster.
In practice, rather than constructing the aforementioned distribution $Q^{\prime}$ explicitly, we can draw samples from $Q^{\prime}$ by sampling from $Q$ and aggregating samples according to their clusters. The modified uncertainty estimating algorithm is shown in Algorithm 2.
Algorithm 2 MI estimator. Python implementation with usage example is given in Appendix A.
1:  Input: $\mu\in{\mathcal{M}}_{1}({\mathcal{X}}^{n})$ any (pseudo-) joint distribution over ${\mathcal{X}}^{n}$ $k\in\mathbb{N}$ sample size $\gamma_{1},\gamma_{2}\geq 0$ stabilization parameters (typically selected as $1/k$) $s:{\mathcal{X}}^{n}\times{\mathcal{X}}^{n}\rightarrow$ a similarity function $\tau\in$ a similarity threshold
3:  Construct a set of indices of unique elements $U=\big{\{}i\in[k]~{}:~{}X_{i}\neq X_{j}\quad\forall j<i\big{\}}$
4:  Construct cluster centers $S\subset U$ according to the similarity function: for all $i,t\in S$, we have $s(X_{i},X_{t})<\tau$ and cluster associated with $X_{i}$ is $D(i)=\big{\{}j\in U~{}:~{}s(X_{i},X_{j})\geq\tau\big{\}}$. Aggregated probabilities: for all $i\in S$,
$\mu^{\prime}(X_{i})=\sum_{j\in D(i)}\mu(X_{j})$
5:  Construct empirical distributions: for all $i\in S$,
$\widehat{\mu}(X_{i})$ $=\frac{\mu^{\prime}(X_{i})}{Z}~{},\quad\text{where}\quad Z=\sum_{j\in S}\mu^{\prime}(X_{j})$
$\widehat{\mu}^{\otimes}(X_{i})$ $=\prod_{j=1}^{n}\sum_{t\in S:X_{t,j}=X_{i,j}}\hat{\mu}(X_{t,1},\ldots,X_{i,j},\ldots,X_{t,n})$
6:  Compute estimate
$\widehat{I}_{k}(\gamma_{1},\gamma_{2})=\sum_{i\in S}\widehat{\mu}(X_{i})\,\ln\left(\frac{\widehat{\mu}(X_{i})+\gamma_{1}}{\widehat{\mu}^{\otimes}(X_{i})+\gamma_{2}}\right)$
The estimator is constructed using only (semantically) equivalent elements in the sample (the indices of these representative elements are collected in $S$), that is, we do not account for duplicate samples and we aggregate probabilities of samples that are lexically different but semantically equivalent. Algorithm 2 works with the aggregated probability distribution $\mu^{\prime}=\widetilde{Q}^{\prime}$ (line 4) by summing over cumulative probabilities over clusters. Note that $D_{\mathrm{KL}}(\mu)\geq D_{\mathrm{KL}}(\mu^{\prime})$ by monotonicity property of KL-divergence (Polyanskiy and Wu, 2024, Theorem 2.16) (this is because $\mu^{\prime}$ is defined on a smaller support). Therefore, Theorem 4.6 implicitly gives a bound on $I(\mu^{\prime})$, and eventually we have $I(\mu)\geq I(\mu^{\prime})$. But we can also directly apply Theorem 4.5 and Theorem 4.6 to distributions $P^{\prime}$ and $Q^{\prime}$ and obtain that $D_{\mathrm{KL}}(\widetilde{Q}^{\prime},\widetilde{P}^{\prime})\geq I(\widetilde{Q}^{\prime})$.
@@ 5 Score-based hallucination tests
Let $\widehat{I}_{k}(\gamma,x)\equiv\widehat{I}_{k}(\gamma)$ computed as in Algorithm 1 for $\mu=\widetilde{Q}$, to emphasize the explicit dependence on the query $x$. The uncertainty estimate $\widehat{I}_{k}(\gamma,x)$ derived above can be used as a score indicating the strength of our belief that the LLM hallucinates for the given query $x$. Such a score can then be used to design abstention policies: if the response is deemed to be hallucinated, the system abstains from responding, while a response is provided otherwise. Score-based abstention methods usually compute a score chosen by the user (such as the response likelihood or the estimator $\widehat{I}(\gamma)$ discussed earlier), and declare hallucination if the score is above or below a threshold, which is determined through calibration. To detect hallucinations successfully, the threshold can be adjusted through calibration on a given task using a hold-out (ground-truth) sample, see, for instance, the paper of Yadkori et al. (2024) where this calibration is discussed in detail.
@@
$a_{\lambda}(x)=\begin{cases}0,&\text{ if }\ \widehat{I}_{k}(\gamma_{1},\gamma_{2},x)<\lambda;\\ 1,&\text{ if }\ \widehat{I}_{k}(\gamma_{1},\gamma_{2},x)\geq\lambda;\end{cases}$
@@ 6 Experiments
In this section we evaluate our abstention method derived based on the MI estimate in Section 5 on a variety of closed-book open-domain question-answering tasks.
@@
Datasets. We consider three different datasets and their combinations: As base datasets, we consider (i) a random subset of $50,000$ datapoints from the TriviaQA dataset (Joshi et al., 2017), and (ii) the entire AmbigQA dataset (with $12038$ datapoints) (Min et al., 2020). These datasets mostly contain single-label queries, and only contain a few multi-label ones. [^7] Note that the multi-label queries in these datasets typically behave as single-label ones in the sense that the LLM assigns overwhelming probability to a dominant response. Moreover, we created a multi-label dataset based on the WordNet dataset (Fellbaum, 1998): We extracted all (6015) datapoints from WordNet at depth $4$ or more of the physical_entity subtree. For each datapoint (entity, children) in WordNet, we constructed a query of the form “Name a type of entity.” and children are considered target labels.
@@
Comparison of responses and computing the output distributions. We use the F1 score888 In this context, the F1 score is calculated based on token inclusion (Joshi et al., 2017, Devlin et al., 2019): for two sequences $a=(a_{1},\ldots,a_{n})$ and $b=(b_{1},\ldots,b_{m})$, defining $p=|a\cap b|/n$ and $r=|a\cap b|/m$ (where $|a\cap b|$ is the size of the intersection of $a$ and $b$, in which for repetitions of an element $y$, we consider the minimum number of repetitions in $a$ and $b$, i.e., $\min_{c\in\{a,b\}}|\{i\,:\,c_{i}=y\}|$, in calculating the size of the intersection) we define $F1=2pr/(p+r)$. Relating to the standard definition of the F1 score, $p$ and $r$ play the role of precision and recall, respectively, if $a$ is thought of as a prediction of $b$. thresholded at $0.25$ to decide if two text sequences match. When multiple responses are sampled, we approximate the output distribution of an LLM in a semantically meaningful way by collapsing matching responses into a single response: we sample $k=10$ responses at temperature $0.9$ for each query, and after eliminating repetitions, all those that match (according to the F1 score) are considered identical and their probabilities are aggregated. We only consider queries for which the greedy (temperature zero) and at least one of the random responses are shorter than $20$ characters. This is because the F1 score (as a match function) and log-probabilities (as a measure of uncertainty) are less reliable for longer sequences. After this filtering, we are left with $38870$ datapoints for TriviaQA, $5315$ datapoints for AmbigQA, and $3296$ datapoints for WordNet.
As shown in prior works (e.g. Kuhn et al. (2023), Yadkori et al. (2024)), we can use LLM self-prompting to obtain more reliable text comparisons specially for longer outputs. Such an approach however is computationally much more expensive.
Algorithm 3 Alternative MI estimator. A usage example is given in Appendix B
1:  Input: $\mu\in{\mathcal{M}}_{1}({\mathcal{X}})$ any distribution over ${\mathcal{X}}$ $k\in\mathbb{N}$ sample size $\gamma_{1},\gamma_{2}\geq 0$ stabilization parameters (typically selected as $1/k$) $s:{\mathcal{X}}\times{\mathcal{X}}\rightarrow$ a similarity function $\tau\in$ a similarity threshold
2:  Independently sample outputs $X_{1},\,\ldots,\,X_{k}\sim\mu\in{\mathcal{M}}_{1}({\mathcal{X}})$
3:  Construct a set of indices of unique elements $U=\big{\{}i\in[k]~{}:~{}X_{i}\neq X_{j}\quad\forall j<i\big{\}}$
4:  Construct cluster centers $S\subset U$ according to the similarity function: for all $i,t\in S$, we have $s(X_{i},X_{t})<\tau$ and cluster associated with $X_{i}$ is $D(i)=\big{\{}j\in U~{}:~{}s(X_{i},X_{j})\geq\tau\big{\}}$. Aggregated probabilities: for all $i,t\in S$,
$\mu^{\prime}_{1}(X_{i})=\sum_{j\in D(i)}\mu(X_{j}),\qquad\mu^{\prime}_{2}(X_{t}\,|\,X_{i})=\sum_{j\in D(t)}\mu(X_{j}\,|\,X_{i})$
5:  Construct empirical distributions: for all $i,t\in S$,
$\widehat{\mu}_{1}(X_{i})$ $=\frac{\mu^{\prime}_{1}(X_{i})}{Z}~{},\quad\text{where}\quad Z=\sum_{j\in S}\mu^{\prime}_{1}(X_{j})$
$\widehat{\mu}_{2}(X_{t}\,|\,X_{i})$ $=\frac{\mu^{\prime}_{2}(X_{t}\,|\,X_{i})}{Z_{i}}~{},\quad\text{where}\quad Z_{i}=\sum_{j\in S}\mu^{\prime}_{2}(X_{j}\,|\,X_{i})$
$\widehat{\mu}(X_{i},X_{t})$ $=\widehat{\mu}_{1}(X_{i})\widehat{\mu}_{2}(X_{t}\,|\,X_{i})\,,\qquad\widehat{\mu}^{\otimes}(X_{i},X_{t})=\widehat{\mu}_{1}(X_{i})\sum_{j\in S}\widehat{\mu}_{1}(X_{j})\widehat{\mu}_{2}(X_{t}\,|\,X_{j})$
6:  Compute estimate
$\widehat{I}_{k}(\gamma_{1},\gamma_{2})=\sum_{i,t\in S}\widehat{\mu}(X_{i},X_{t})\,\ln\left(\frac{\widehat{\mu}(X_{i},X_{t})+\gamma_{1}}{\widehat{\mu}^{\otimes}(X_{i},X_{t})+\gamma_{2}}\right)$
@@
Baselines. We consider abstention policies based on four scoring methods. The first three are as follows: (i) the probability of the greedy response (denoted by $T0$); (ii) the semantic-entropy method of Kuhn et al. (2023) whose score is the entropy of $k=10$ generated samples (denoted by S.E.). To calculate entropy, we first aggregate probabilities of equivalent responses and normalize the probabilities so that they sum to 1 (as described above); and (iii) our proposed mutual information score as defined in Section 4 (and denoted by M.I.) with the choices of $k=10$, $n=2$, and $\gamma_{1}=\gamma_{2}=0$ (the latter choice approximates the case that the number of potential responses can be very large in which case the theoretical choice of $\gamma_{1}$ and $\gamma_{2}$ would be very small). To calculate the mutual information, as shown in Algorithm 3, we first generate $k=10$ random samples. Then for any response $Y$, we calculate the probability of all generated responses given the prompt $F_{1}(x,Y)$. We construct estimates $\widehat{Q}(Y)$ and $\widehat{Q}(Y^{\prime}|Y)$ by aggregating probabilities of equivalent responses, and normalizing the probabilities so that they sum to 1.
The calculation of the mutual information is slightly different than the algorithms presented in Algorithm 1 and Algorithm 2 and takes advantage of the available log-likelihood function in LLMs. Notice that the input $\mu$ in Algorithm 3 is LLM distribution $Q$ as opposed to being the pseudo joint distribution $\widetilde{Q}$ in Algorithm 1. Another difference is that the similarity function $s$ now takes two texts as input (as opposed to taking two $n$-dimensional arrays of texts as inputs in Algorithm 2). As explained earlier, we use the F1 score as the similarity function and we use $\tau=0.25$ as the similarity threshold.
@@
In our experiments we either sweep through all abstention thresholds (Figure 5), or optimize the threshold on some calibration data, as explained in the description of the relevant experiment (Figure 6).
Results. We consider the precision-recall (PR) trade-off for the various methods on the different datasets. Here, recall is the percentage of queries where the method does not abstain, and precision is the percentage of correct decisions among these queries. [^10] In some figures, for better illustration, we show the error rate which is one minus the precision. Figure 5ab show PR-curves for the baselines and the proposed method on TriviaQA and AmbigQA. As can be seen, our method is better than the $T0$ and S.V. baselines, but performs similarly to the S.E. method. This is because the TriviaQA and AmbigQA datasets contain mostly single-label queries, and therefore a first-order method such as S.E. is sufficient to detect hallucinations. The AmbigQA dataset contains a few multi-label queries, but upon closer inspection, we observe that the LLM has low entropy on most of these queries. [^11] Such a case can also be seen in the query “Name a city in the UK.” in Figure 3 where the response “London” has probability $0.958$. Therefore, a first-order method can perform as well as our method on such queries. Our proposed method, as well as the baselines, make no mistakes on the WordNet dataset (as the prediction of the LLM is always correct), hence we omit those results. The S.V. baseline performs significantly worse than the other methods when the recall is not high (is below about 0.8).
@@
The similar performance for the S.E. and M.I. methods shown in Figure 5ab is due to the fact that the LLM has low entropy on most multi-label queries. However, ideally, an LLM should have higher entropy on multi-label queries (which would demonstrate broader knowledge, not focusing on a single possible answer). To include such queries, we mix the TriviaQA and AmbigQA datasets with our WordNet-based dataset with “truely” multi-label queries as constructed above. To enhance the intended effect, we filter our WordNet dataset by keeping only queries with entropy higher than $0.7$ (approximately the entropy of the uniform distribution over two atoms). Then we have $842$ remaining datapoints in WordNet. Note that when considered in isolation, both our proposed method and the semantic entropy method rarely make mistakes on this dataset. Then we create two new datasets by combining our $842$ WordNet datapoints with $842$ randomly selected datapoints from TriviaQA and AmbigQA, respectively, resulting in the TriviaQA+WordNet and AmbigQA+WordNet datasets. Figure 5cd show PR-curves for the S.E. and M.I. methods on these two combined datasets. Apart from low recall values, the performance of the S.E. method degrades noticeably with the addition of extra multi-label data. This precision/recall curve might look somewhat strange (with precision sometimes increasing with recall); this is due to the fact that both methods are always correct on the large number of high-entropy WordNet queries, where the LLM’s default predictions are correct.
@@
The hardness with the combined datasets is that the predominantly single-label datasets (TriviaQA, AmbigQA) might need a different calibration threshold than the multi-label WordNet dataset, and this is better handled by our proposed method than by S.E. To better illustrate the improved abstention properties of our method, we examine how the two methods handle when the output of the LLM is diverse (i.e., has high entropy). In order to do this, we perform the following experiment: We create a calibration dataset by adding $500$ random datapoints from the WordNet dataset to $500$ random datapoints from TriviaQA, and another such random dataset for test. We determine the abstention thresholds on the calibration dataset for both the S.E. and the M.E. methods, [^12] This is done by fixing the target loss rates of 0.05 for TriviaQA and 0.15 for AmbigQA, and finding threshold parameters that lead to these rates on the calibration set. and measure the performance (error rate, i.e., 1 minus precision, and recall) of the resulting abstention policies on the test set. We repeat this process $10$ times and report mean values and 95% confidence intervals with Gaussian approximation. We perform a similar evaluation process for mixtures of AmbigQA and WordNet datasets. Figure 6 show that while the S.E. method has similar recall and error rates to those of the proposed method on low-entropy queries, its recall values are much lower for queries with higher entropy, while the M.E. method makes only few mistakes on these queries.
@@
Figure 6: Recall and error rates (one minus precision: percentage of mistakes when not abstaining) of the proposed and the baseline method on TriviaQA+WordNet and AmbigQA+WordNet datasets. On TriviaQA+WordNet and AmbigQA+WordNet datasets, the methods are calibrated at target loss of $0.05$ and $0.15$, respectively. On the x-axis, the queries are partitioned according to the entropy of the LLM’s output. Error bars show 2 standard deviation confidence intervals (based on 10 repetitions). While the first-order S.E. method has similar recall and error rates to those of the proposed M.E. method on low-entropy queries, its recall values are nearly zero for queries with higher entropy.